OpenStudy (beccad97):

Evan is landscaping his backyard. The yard is shaped like a rectangle and measures 80 feet by 70 feet. He wants to spread topsoil evenly over the entire surface. One load of topsoil will cover 400 square feet, 4 inches deep. How many loads of dirt does Evan need in order to cover his entire yard?

OpenStudy (mertsj):

Use unit analysis: \[5600 ft^2 \times \frac{1 load}{400 ft^2}= \frac{5600 load}{400}\]

OpenStudy (mertsj):

You can finish the simplification.